# LLM4ED: Large Language Models for Automatic Equation Discovery

### v1

@@ 3.4.1 Constant optimization
For ODEs, constants can appear at any position in the expression tree. We first generate equation structures through LLMs, namely "skeleton", and then utilize the Broyden-Fletcher-Goldfarb-Shanno algorithm (BFGS) [50] to execute the following optimization objective.
@@
Rounds of optimization iterations are performed using Scipy.Minimize to ultimately determine all the constants in the expression tree. Note that if the generated equations do not contain the constant operator, sparse regression techniques can be employed to assign coefficients that are not 1 to each term, thereby improving the accuracy of the discovered equations and facilitating the identification of lengthy true equations.
Figure 3: Workflow of the proposed framework.
@@ 3.5 Optimization
This study utilizes two LLM-guided optimization techniques to enhance the optimization process. The self-improvement method primarily performs local modifications based on the equation’s performance, while the genetic algorithm-based approach is conducted for a global search on the elite equations. Our goal is to achieve a better balance between exploration and exploitation.
@@ 3.5.1 Self-improvement process
LLMs have been demonstrated in numerous experiments to function as gradient-free optimizers, possessing the ability to draw inferences from historical data and iteratively optimize to produce superior samples [34]. we include historical elite equations and their corresponding scores as equation-score pairs within the prompt, enabling LLMs to perform local modifications using these data. The modifications primarily encompass two facets: (1) recognizing and eliminating redundant equation terms by leveraging historical data; (2) incorporating and generating novel random equation terms built upon existing equations. These two operations resemble the introduction of "delete" and "add" operators, which can effectively utilize the historical elite samples and aptly supplement the unstable updating of genetic algorithms. An example case of the self-improvement process is shown in Fig. 4 and the customized prompts are demonstrated in Appendix A.
@@ 3.5.2 Equation evolution process
Genetic algorithms are one of the commonly used global optimization methods inspired by natural selection [51, 52]. Evolutionary operators can be applied to the parent individuals to generate new offersprings. In particular, this procedure required an intricate design and application on tree structures in symbolic regression. In this paper, we employ natural language to guide LLMs in the execution of the genetic algorithms, rather than relying on manual coding. Specifically, we conduct crossover and mutation operations on the $M$ equation populations generated in the past, thereby producing a greater variety of equation combinations. The process consists of two steps:
Select parent population
Historical elite equations will be incorporated into the prompt for the evolution process, originating from two sources: a predefined priority queue caching the top $K$ historically elite equations and high-quality samples selected from the last iterations. By combining them, we ultimately retain the $M$ better-performing equations as the parent population.
@@
On one hand, As optimization in genetic algorithms often involves substantial randomness, despite having strong global search abilities, the efficiency of optimization remains suboptimal. Consequently,
3.6 Prompt Engineering
Throughout the entire equation discovery process, our generation and optimization rely on natural language-based prompts, which follow a similar structure in the three processes of initialization, evolution, and self-improvement. The standard format consists of the following components, as shown in Fig. 1.
  • Task descriptions: This part primarily explains the main task and defines the symbol library including operators (e.g., +,+,-+ , -… and operands (e.g., $x,const$).
  • Historical examples: We pass $M$ high-quality equations generated in the past as historical information to LLMs. A priority queue of the Top $K$ expressions in history is established to ensure stable and efficient optimization. All expressions within the priority queue are incorporated into the prompts. Furthermore, we select $M-K$ expressions from the last iteration to preserve sampling diversity, as illustrated in Fig. 2. The manifestation of these samples within the prompt varies according to the optimization techniques employed. In the evolution process, only high-quality equations with string format are presented, while in the self-improvement process, historical samples are shown in the form of equation-score pairs.
  • Instructions: This part is aimed at guiding LLMs to optimize and generate new equations. In the initialization stage, LLMs are primarily required to freely combine symbols from the library to produce equations of arbitrary form and length. In the optimization stage, LLM is mainly guided to generate optimized equations based on the form of historical equations, using different evolutionary operators or making local modifications (i.e. self-improvement).
  • Other hints or constraints: In the initialization stage, we can directly impose requirements on the form of the generated equations through natural language, acting as constraints to limit the search space. In the optimization stage, we can further define local modifications and evolutionary operators, and provide possible examples as a few-shot prompt.
The utilized prompts in this paper are demonstrated in A.
@@ 4.1 Evaluation metrics
where $n$ denotes the total number of function terms; $\xi_{i}$, $\xi_{i}^{*}$ refer to the true coefficients and identified coefficients, respectively. ODEs are more complex in symbolic form. A skeleton with defined symbols need to be constructed first and followed by the optimization of constants within the skeleton, which may generate more symbol combinations. Compared to identifying the most consistent expression in symbolic form, it is more crucial and meaningful to conduct numerical evaluation. Specifically, we aim to find an effective $\hat{\mathcal{F}}$, whose solution trajectories approximate the observed $x$ in the current numerical domain, i.e., all of the expressions are evaluated by the reconstruction accuracy. Furthermore, the other critical criterion is that the solution of the identified $\hat{\mathcal{F}}$ precisely fits the correct trajectories even when the initial condition varies. We utilize the coefficient of determination ($R^{2}$) as the metric to evaluate the agreement between the solution trajectories and the true trajectories: $R^{2}=1-\frac{\sum_{i}^{n}\left(x_{i}-\hat{x}_{i}\right)^{2}}{\sum_{i}^{n}\left(x_{i}-\bar{x}\right)^{2}}\in(-\infty,1]$, where $x_{i}$ denote observations and $\hat{x}_{i}$ refer to predicted values.
@@ 4.2 Experiment settings
The hyperparameters used in the experiments are shown in Table 1. In terms of the experimental setup, the symbol library and equation assumptions used for mining ODE and PDE equations are slightly different, as shown in Table 2. The library used for mining PDEs has relatively fewer operators and more operands involved, and does not include the symbol "const". On the other hand, the library used for mining ODEs covers more mathematical operators and the constants are determined using nonlinear optimization methods, e.g., BFGS [50].
@@
$M$ 20 Number of expressions generated at each iteration
@@
$tol$ $1\times 10^{-4}$ Threshold of weights for reserved function terms
@@ 4.3.1 Equations and discovered results
Table 3: Summary of several nonlinear systems governed by PDEs and discovered results. The subscripts $m$, $n$ denote the number of discretizations.
@@
In the experiments of PDE discovery, we testify the framework’s ability to discover the governing equations of six canonical nonlinear systems, including Burges’ equation, Chafee-Infante equation, PDE_divide equation with fractional structure, Kuramoto-Sivashinsky equation with fourth-order derivatives, nonlinear Fisher-KPP equations with a square of the spatial derivative, and two-dimensional Navier-Stokes equation. With the default parameter configuration, our approach is capable of accurately identifying the correct structure of the equations while preserving a minimal coefficient error, as shown in Fig. 3. Notably, in comparison to fixed candidate set methods, our framework diminishes the dependence on prior knowledge, enabling the discovery of more complex equation forms, for example, equations with fractional or compound structures.
4.3.2 Compariosn of different optimization strategies
@@
We further verified the effectiveness of the proposed LLM-guided iterative optimization. Three optimization methods are mainly discussed and compared: (1) only the self-improvement optimization method utilized; (2) only the genetic algorithm utilized; (3) the alternating iterative method of the above two methods (proposed in the framework). The identification experiments for the aforementioned equations were replicated ten times, with each experiment’s maximum iteration count set to 50 iterations, to enable further examination of the performance of various methods. Fig. 6 illustrates that the iterative approach combining both methods yielded the highest frequency of discovering the correct equations, with recovery rates consistently surpassing 80%, outperforming the outcomes achieved by employing a single optimization technique. Fig. 6(b) depicts the successful rate of the ultimately identified equations. It is noteworthy that despite the self-improvement method outperforming the genetic algorithm and exhibiting higher optimization efficiency in some systems, such as Burgers’ equation, it is more prone to converging to local optima. When the iteration step is extended to 100 steps, the symbolic success rate of optimization employing the genetic algorithm approach surpasses 80% for all of the equations, demonstrating its superior global optimization capability, whereas self-improvement hardly achieves significant improvement.
@@
We provide further detailed explanations with the Chafee-Infante, Burgers, and NS equations utilized as examples. Fig. 7 illustrates the evolution of the highest score up to the current iteration step. It reveals that the optimization efficiency of the alternating approach combining both methods is superior, facilitating faster identification of the correct equations in various equation discovery tasks. Utilizing the Chafee-Infante equation as an illustration, we further examine the density distribution of scores at each iteration throughout the optimization process. Fig. 8 illustrates that the self-improvement strategy exhibits a propensity for local modifications on historical elite equations, resulting in a score distribution resembling an incremental trend. Conversely, the GA method leans towards global searches, identifying equations with higher diversity, albeit at the cost of potentially compromised optimization efficiency. Employing alternating iterations of both methods proves to be more advantageous in striking a balance between exploration and exploitation.
@@
Figure 7: Evolution of the best score with different methods while discovering the Chafee-Infante equation, Burgers’ equation and NS equation.
@@
Figure 8: Distribution of scores at different iterations with different methods while discovering the Chafee-Infante equation.
@@ 4.4 ODE discovery task
ODE-12 $c_{0}\sin{\left(x\right)}^{2}+c_{1}x+c_{2}\cos{\left(x\right)}+c_{3}$ [0.464,0.907,2.7834,-2.7836] 0.999 0.999
@@
In this section, we testified our framework on 16 one-dimensional ODEs in a comprehensive benchmark named ODE-bench [39], which has been utilized to describe real-world phenomena by Strogatz [53]. The equation information is listed in the Appendix B. Each equation contains two sets of trajectories with two different initial conditions. We take one set of trajectory data as training data and search to find the optimal $\hat{\mathcal{F}}$. During the evaluation process, we consider the solution trajectory of the ODE equation associated with $\mathcal{F}$ as predicted results and utilize $R^{2}$ score as the evaluation criterion to measure the fitting accuracy in comparison to the actual trajectory. The value $R^{2}$ on the training set represents the accuracy of the reconstruction, while the value $R^{2}$ on the testing set with a new initial condition signifies the generalization performance. As shown in Table 4, we repeated the search process for each equation three times and provided the best results among them. It can be seen that in our framework, the percentage of equations with $R^{2}$ greater than 0.99 on the training set is 93.75% (15/16), and on the test set, the percentage of equations with R2 greater than 0.99 is 68.75%. Equations with $R^{2}$ greater than 0.9 exceeded 90% on both the testing and training sets. Fig. 9 presents the detailed prediction results for each equation.
Figure 9: Predicted solution trajectories on training set and testing set.
Table 5: Evaluation on 16 ODEs with different methods and LLMs. We counted the number of equations discovered by different methods and LLMs that meet the corresponding $R^{2}$ criteria. When the $R^{2}$ value is negative, or the solution trajectories of the ODE associated with $\hat{\mathcal{F}}$ exhibit numerical overflow, we deem the discovered equation as "Invalid".
@@
GPT-3.5-turbo 13 (81.25%) 14 (87.5%) 3 (18.75%) 12 (75%) 13 (81.25%) 3 (18.75%) 8 (50%)
GPT-4 15 (93.75%) 16 (100%) 0 (0%) 11 (68.75%) 15 (93.75%) 1 (6.25%) 8 (50%)
@@
We conducted a further comparison of the performance of different equation discovery methods and large models on all the presented ODEs. Three symbolic regression methods are utilized as the baseline models, including PySR [54] and ODEformer [39]. PySR is a practical and high-performance library for symbolic regression based on a multi-population evolutionary algorithm. PySR is aimed at single-instance datasets and has been broadly adopted for interpretable symbolic discovery. ODEformer is based on pretrained transformers and achieves SOTA on ODEbench’s datasets. The above two methods are experimented with according to their default hyperparameter configurations.
@@
As illustrated in Table 4, our framework demonstrates equivalent reconstruction performance to the above methods designed for symbolic regression, while exhibiting superior generalization capabilities and usability. Note that PySR, which is based on evolutionary search, possesses powerful search capabilities and can discover equations that are numerically accurate on the training dataset. However, these discovered equations tend to have higher complexity and consequently are susceptible to exhibiting poor generalization performance. Additionally, we performed a comparative analysis of the performance of large models with varying parameter sizes and language capacities on this task. We take the open-source large language model Llama2 with 7B parameters  [55] and more advanced models, including GPT-3.5-turbo and GPT-4. The results indicate that as the capability of the large model improves, the identified equations become relatively more accurate both on the training set and testing set. This can be primarily attributed to the fact that the capacity of the large model directly exerts a substantial influence on the generation and optimization of equations. On one hand, large models with inferior capabilities may struggle to effectively understand and execute the provided instructions, such as the constraints we defined, leading to the generation of numerous invalid equations. Furthermore, they tend to fail to properly execute GA instructions and fail to accurately perform crossover and mutation operations. Conversely, the model’s reasoning capacity directly influences its self-improvement optimization capabilities. It is also worth noting that as the capabilities and number of parameters of large models further increase, the gains in accuracy are diminishing, especially on test sets.
@@
In addition, we need to further emphasize the efficiency of the proposed framework. The total running time is the product of the time per iteration and the total number of iterations. The time cost per iteration ranges from 10s to 40s under the default configuration and mainly includes the time spent on remotely accessing the LLM API and evaluating the feedback of LLM, i.e., the generated equations. The time consumed for accessing the API interface exhibits a positive correlation with the number of samples generated and is roughly order of magnitude larger than the time for evaluations. In practice, we can allocate separate processes to query the LLM in parallel. On one hand, this approach can reduce the total response time. On the other hand, we can increase the number of expressions generated in each iteration, which in turn contributes to more accurate samples for optimization and helps reduce the number of optimization iterations.

### v2

@@ 3.5.1 Constant optimization
For ODEs, constants can appear at any position in the expression tree. We first generate equation skeleton through LLMs, and then utilize the Broyden-Fletcher-Goldfarb-Shanno algorithm (BFGS) [52] to execute the following optimization objective.
@@
Rounds of optimization iterations are performed using scipy.optimize.minimize to ultimately determine all the constants in the expression tree. Note that if the generated equations do not contain constant operands, sparse regression techniques can be employed to assign coefficients other than 1 to each term, thereby improving the accuracy of the discovered equations and facilitating the identification of lengthy true equations.
Figure 3: Determination of PDE coefficients.
@@ 3.6 Optimization
This study utilizes two LLM-guided optimization techniques to enhance the equation refinement process. The self-improvement method primarily performs local modifications based on the equation’s performance, while the genetic algorithm-based approach is employed for a global search on the elite equations. Our goal is to achieve a better balance between exploration and exploitation.
@@ 3.6.1 Self-improvement process
LLMs have been demonstrated in numerous experiments to function as gradient-free optimizers, possessing the ability to draw inferences from historical data and iteratively optimize to produce superior samples [34]. We include historical elite equations and their corresponding scores as equation-score pairs within the prompt, enabling LLMs to perform local modifications using these data. The modifications primarily encompass two facets: (1) recognizing and eliminating redundant equation terms by leveraging historical data; (2) incorporating and generating novel random equation terms built upon existing equations. These two operations resemble the introduction of "delete" and "add" operators, which can effectively utilize the historical elite samples and aptly supplement the potentially unstable updating of genetic algorithms. An example case of the self-improvement process is shown in Fig. 4 and the customized prompts are demonstrated in Appendix A.
@@ 3.6.2 Equation evolution process
Genetic algorithms are one of the commonly used global optimization methods inspired by natural selection [53, 54]. Evolutionary operators can be applied to the parent individuals to generate new offspring. In particular, this procedure requires an intricate design and application on tree structures in symbolic regression. In this paper, we employ natural language to guide LLMs in the execution of the genetic algorithms, rather than relying on manual coding. Specifically, we conduct crossover and mutation operations on the $M$ equation populations generated in the past, thereby producing a greater variety of equation combinations. The process consists of two steps:
Construct parent population
Historical elite equations will be incorporated into the prompt for the evolution process, originating from two sources: a predefined priority queue caching the top $K$ historically elite equations and high-quality samples selected from the last iteration. By combining them, we ultimately retain the $M$ better-performing equations as the parent population.
@@
Self-improvement based on local modifications effectively utilizes the reasoning capabilities of LLMs to identify the direct mapping relationships between symbol combinations and scores, thereby refining equations in a gradient-free manner. However, this approach is prone to becoming trapped in local optima. On the other hand, genetic algorithms exhibit strong global optimization abilities, but their effectiveness heavily depends on the quality of the initial population, resulting in possible updating instability. By integrating these two strategies and employing a dynamic buffer of high-quality equations, the evolution and refinement process of equations can be significantly complemented and enhanced.
@@ 4.1 Evaluation metrics
where $n$ denotes the total number of function terms; $\xi_{i}$, $\xi_{i}^{*}$ refer to the true coefficients and identified coefficients, respectively. ODEs are more complex in symbolic form. A skeleton with defined symbols needs to be constructed first, followed by the optimization of constants within the skeleton, which may generate more symbol combinations. Compared to identifying the most consistent expression in symbolic form, it is more crucial and meaningful to conduct numerical evaluation. Specifically, we aim to find an effective $\hat{\mathcal{F}}$, whose solution trajectories approximate the observed $x$ in the current numerical domain, i.e., all of the expressions are evaluated by the reconstruction accuracy. Furthermore, the other critical criterion is that the solution of the identified $\hat{\mathcal{F}}$ precisely fits the correct trajectories even when the initial condition varies. We utilize the coefficient of determination ($R^{2}$) as the metric to evaluate the agreement between the solution trajectories and the true trajectories: $R^{2}=1-\frac{\sum_{i}^{n}\left(x_{i}-\hat{x}_{i}\right)^{2}}{\sum_{i}^{n}\left(x_{i}-\bar{x}\right)^{2}}\in(-\infty,1]$, where $x_{i}$ denotes observations and $\hat{x}_{i}$ refers to predicted values.
@@ 4.2 Experiment settings
The hyperparameters used in the experiments are shown in Table 1. GPT-3.5-turbo is utilized as the default LLM backbone. In terms of the experimental setup, the symbol library and equation assumptions used for mining ODEs and PDEs are slightly different, as shown in Table 2. The library used for mining PDEs has relatively fewer operators and more operands involved and does not include the symbol "const". On the other hand, the library used for mining ODEs covers more mathematical operators, and constants are determined using nonlinear optimization methods, e.g., BFGS [52].
@@
$M$ 10 Number of expressions generated at each iteration
@@
$T$ 0.9 LLM temperature
@@ 4.3.1 Equations and discovered results
Table 3: Summary of canonical nonlinear systems governed by PDEs and discovered results. The subscripts $m$ and $n$ denote the number of discretizations.
@@
In the experiments of PDE discovery, we demonstrate the framework’s ability to discover the governing equations of six canonical nonlinear systems, including the Burges’ equation, Chafee-Infante equation, PDE_divide equation with fractional structure, Kuramoto-Sivashinsky (KS) equation with fourth-order derivatives, nonlinear Fisher–Kolmogorov–Petrovsky–Piskunov (Fisher-KPP) equations with a square of the spatial derivative, and two-dimensional Navier-Stokes (NS) equation. With the default parameter configuration, our approach accurately identifies the correct structure of the equations while maintaining minimal coefficient error, as shown in Fig. 3. Notably, in comparison to fixed candidate set methods, our framework reduces the dependence on prior knowledge, enabling the discovery of more complex equation forms, such as equations with fractional or compound structures.
4.3.2 Comparison of different optimization strategies
@@
We further verified the effectiveness of the proposed LLM-guided iterative optimization. Three optimization methods are primarily discussed and compared: (1) using only the self-improvement optimization method; (2) using only the genetic algorithm; and (3) the alternating iterative method combining the two methods (as proposed in the framework). The identification experiments for the aforementioned equations were replicated ten times, with each experiment’s maximum iteration count set to 50, to further examine the efficiency of various methods. Fig. 6 illustrates that the iterative approach combining both methods yielded the highest frequency of discovering the correct equations, with recovery rates consistently surpassing 80%, outperforming the outcomes achieved by employing a single optimization technique. Fig. 6(b) depicts the success rate of the ultimately identified equations. It is noteworthy that despite the self-improvement method outperforming the genetic algorithm and exhibiting higher optimization efficiency in some systems, such as the Burgers’ equation, it is more prone to converging to local optima. When the iteration count is extended to 100 steps, the symbolic success rate of optimization employing the genetic algorithm approach surpasses 80% for all of the equations, demonstrating its superior global optimization capability, whereas self-improvement hardly achieves significant improvement.
@@
We provide further detailed analysis with the Chafee-Infante, Burgers, and NS equations utilized as examples. Fig. 7 illustrates the evolution of the maximum score throughout the optimization process. It reveals that the optimization efficiency of the alternating approach combining both methods is superior, facilitating faster identification of the correct equations in various equation discovery tasks. Utilizing the Chafee-Infante equation as an illustration, we further examine the density distribution of scores at each iteration throughout the optimization process. Fig. 8 illustrates that the self-improvement strategy exhibits a propensity for local modifications on historical elite equations, resulting in the overall score resembling an incremental trend. Conversely, GA excel at global searches, identifying equations with higher diversity, albeit at the cost of potentially compromised optimization efficiency. Employing alternating iterations of both methods proves more advantageous in striking a balance between exploration and exploitation.
@@
Figure 7: Evolution of the maximum score with different methods while discovering the Chafee-Infante equation, Burgers’ equation and NS equation.
@@
Figure 8: Distribution of scores at different iterations with different methods while discovering the Chafee-Infante equation. Numbers in the figure refer to different iteration steps.
@@ 4.4.1 Discovered results
ODE-12 $c_{0}\sin^{2}{\left(x\right)}+c_{1}x+c_{2}\cos{\left(x\right)}+c_{3}$ [0.464,0.907,2.7834,-2.7836] 0.999 0.999
@@
In this section, we tested our framework on 16 one-dimensional ODEs in a comprehensive benchmark named ODE-bench [41], which has been utilized to describe real-world phenomena by Strogatz [55]. The equation information is listed in Appendix B. Each equation contains two sets of trajectories with two different initial conditions. We used one set of trajectory data as training data and searched to find the optimal $\hat{\mathcal{F}}$. During the evaluation process, we consider the solution trajectory of the ODE associated with $\mathcal{F}$ as predicted results and utilize the $R^{2}$ score as the evaluation criterion to measure the fitting accuracy in comparison to the actual trajectory. The $R^{2}$ value on the training set represents the accuracy of the reconstruction, while the $R^{2}$ value on the test set with a new initial condition signifies the generalization performance. As shown in Table 4, we repeated the search process for each equation three times and reported the best results among them. It can be seen that in our framework, the percentage of equations with $R^{2}$ greater than 0.99 on the training set is 93.75% (15/16), and on the test set, the percentage of equations with $R^{2}$ greater than 0.99 is 68.75%. Equations with $R^{2}$ greater than 0.9 exceeded 90% on both the training and test sets. Fig. 9 presents detailed prediction results for each equation. Most solution trajectories are consistent with the true trajectories.
Figure 9: Predicted solution trajectories on training and test sets.
Table 5: Evaluation on 16 ODEs with different methods and LLM backbones. We counted the number of equations discovered by different methods and LLMs that meet the corresponding $R^{2}$ criteria. When the $R^{2}$ value is negative, or the solution trajectories of the ODE associated with $\hat{\mathcal{F}}$ exhibit numerical overflow, we deem the discovered equation as "Invalid".
@@
GPT-3.5-turbo 13 (81.25%) 13 (81.25%) 3 (18.75%) 12 (75%) 13 (81.25%) 3 (18.75%) 8 (50%)
GPT-4 15 (93.75%) 16 (100%) 0 (0%) 11 (68.75%) 15 (93.75%) 0 (0%) 8 (50%)
4.4.2 Comparisons with SR benchmarks and ablation studies
@@
We conducted a further comparison of the performance of SR benchmarks. Two symbolic regression benchmarks are utilized as the baseline models, including PySR [56] and ODEformer [41]. PySR is a practical and high-performance library for symbolic regression, based on a multi-population evolutionary algorithm. PySR is designed for single-instance datasets and has been broadly adopted for interpretable symbolic discovery. ODEformer is based on pretrained transformers and achieves SOTA on ODEbench’s datasets. The aforementioned methods were tested according to their default hyperparameter configurations. PySR is implemented with a population size of 40 (the number of populations running) and total iterations of 40. ODEformer is implemented with a beam size of 20 and a temperature of 0.1.
As illustrated in Table 4, our framework demonstrates equivalent reconstruction performance to the aforementioned methods designed for symbolic regression, while exhibiting superior generalization capabilities and usability. Note that PySR, which is based on evolutionary search, possesses powerful search capabilities and can discover equations that are numerically accurate on the training dataset. However, these discovered equations tend to have more complex structures and, consequently, are susceptible to exhibiting poor generalization performance.
@@
Additionally, we performed a comparative analysis of the performance of LLM backbones with varying parameter sizes and language capabilities on this task. We used the open-source large language model Llama2 with 7B parameters [57] and more advanced models, including GPT-3.5-turbo and GPT-4. The results indicate that as the capability of LLMs improves, the identified equations become relatively more accurate on both the training set and the test set. This can be primarily attributed to the fact that the capability of the LLM directly exerts a substantial influence on the generation and optimization of equations. On the one hand, large models with inferior capabilities may struggle to effectively understand and execute the provided instructions, such as the constraints we defined, leading to the generation of numerous invalid equations. Furthermore, they tend to fail to properly execute GA instructions and accurately perform crossover and mutation operations. Conversely, the model’s reasoning capability directly influences its self-improvement optimization capabilities. It is also worth noting that as the capabilities and number of parameters of large models further increase, the gains in accuracy diminish, especially on test sets.
@@
In addition, we further emphasize the efficiency of the proposed framework. The total running time is the product of the time per iteration and the total number of iterations. The time cost per iteration ranges from 10s to 40s under the default configuration and mainly includes the time spent on remotely accessing the LLM API and evaluating the feedback from the LLM, i.e., the generated equations. The time consumed for accessing the API interface exhibits a positive correlation with the number of samples generated and is roughly an order of magnitude larger than the time for evaluations. In practice, we can allocate separate processes to query the LLM in parallel. On one hand, this approach can reduce the total response time. On the other hand, we can increase the number of expressions generated in each iteration, which in turn contributes to more accurate samples for optimization and helps reduce the number of optimization iterations.